### v2

# FSPGD: Rethinking Black-box Attacks on Semantic Segmentation

Abstract
Transferability, the ability of adversarial examples crafted for one model to deceive other models, is crucial for black-box attacks. Despite advancements in attack methods for semantic segmentation, transferability remains limited, reducing their effectiveness in real-world applications. To address this, we introduce the Feature Similarity Projected Gradient Descent (FSPGD) attack, a novel black-box approach that enhances both attack performance and transferability. Unlike conventional segmentation attacks that rely on output predictions for gradient calculation, FSPGD computes gradients from intermediate layer features. Specifically, our method introduces a loss function that targets local information by comparing features between clean images and adversarial examples, while also disrupting contextual information by accounting for spatial relationships between objects. Experiments on Pascal VOC 2012 and Cityscapes datasets demonstrate that FSPGD achieves superior transferability and attack performance, establishing a new state-of-the-art benchmark. Code is available at https://anonymous.4open.science/r/FSPGD/README.md.
*Corresponding author
1 Introduction
Convolutional neural networks (CNNs) have shown remarkable capabilities across a range of domains, including image classification [21, 45, 23, 46], semantic segmentation [6, 7, 58, 34], and image synthesis [39, 14, 40, 41, 42], and have consistently achieved state-of-the-art performance. However, their vulnerability to adversarial attacks, which are strategically crafted perturbations that lead to misclassification or incorrect predictions, remains a significant concern. The presence of such vulnerabilities raises some issues, particularly in security-sensitive applications like autonomous driving [13] and facial verification [44]. To address this problem, various adversarial attack methods have been studied [48, 31, 52, 4, 57, 17, 28, 36, 11, 32, 54, 10, 19, 25, 50, 56], but it has not yet been fully resolved.
Adversarial attacks are categorized as white-box and black-box attacks [48, 52]. In a white-box attack, the attacker has complete knowledge of the target model, including its architecture, parameters, and gradients, enabling precise crafting of adversarial examples. While white-box attacks show strong attack performance, they often exhibit lower transferability, limiting their effectiveness in real-world applications [48, 11, 54, 19]. Conversely, black-box attacks assume no prior knowledge of the model structure or parameters. Instead, the attacker relies on querying the model and analyzing outputs to generate adversarial examples. Although more challenging, black-box attacks are more suitable for real-world applications where model specifics are unknown. This paper aims to analyze limitations in existing black-box attack methods and introduce a novel approach to address these challenges.
In the black-box attack, the ability of adversarial examples generated for source model to deceive target models, which is called transferability, is a crucial property. However, enhancing the transferability is challenging since different CNN models learn and represent distinct features. This variation makes it difficult for adversarial examples generated for a source model to generalize effectively to target models. To resolve this problem, various black-box attack methods, such as data [36, 11, 32, 54, 49], optimization [10, 32, 19, 35], feature [25, 50, 56, 30, 51] and model [59, 29, 18, 60] perspectives, have been explored in the field of image classification. Although these methods show strong attack performance and transferability in image classification tasks, applying them directly to semantic segmentation, which requires classifying each pixel in the input image, is challenging.
To overcome this problem, various adversarial attack methods [17, 26, 53, 1, 5, 4, 27] specifically designed for semantic segmentation have been introduced. While these methods show fine attack performance in semantic segmentation, they have not yet fully overcome the challenges of transferability. In this study, we analyze the reasons for the weak transferability of existing methods and identify the following causes: conventional methods usually calculate gradients and generate perturbations by using the output predictions of the source model. This approach exhibits strong attack performance only on the source model but fail to achieve similar performance on new target models. This limitation arises because these methods only consider pixel-wise predictions and do not effectively attack contextual information, i.e. the spatial relationships between objects, which is a critical factor in semantic segmentation.
To address this problem, this paper proposes a novel black-box attack method, called the Feature Similarity Projected Gradient Descent (FSPGD) attack, which demonstrates strong attack performance and significant transferability. Unlike existing segmentation attack methods that rely solely on output predictions from the source model to compute gradients, the proposed method calculates gradients by leveraging features extracted from the intermediate layer. Specifically, we develop a novel loss function that targets local information by comparing features between clean images and adversarial examples, while also disrupting contextual information by leveraging spatial relationships between objects within the image. To validate the superiority of the proposed method, we present extensive experimental results across a variety of models, such as PSPNet-ResNet50 [58], DeepLabv3-ResNet50 [6], SegFormer-MiT B0 [55], and Mask2Former-Swin S [8]. Moreover, a series of ablation studies are conducted to highlight the robust generalization capabilities of the proposed method. Quantitative evaluations clearly show that the proposed method not only achieves strong attack performance but also surpasses conventional methods in transferability, setting a new state-of-the-art benchmark. Our contribution can be summarized as follows:
  • We investigate the causes of weak transferability in existing segmentation attack methods and propose a novel method, called FSPGD, to address this issue.
  • This paper is the first to apply intermediate feature attacks to the field of semantic segmentation. Through various experiments, we prove that intermediate feature attacks are effective not only in image classification but also in semantic segmentation.
  • We perform extensive experiments on multiple baseline models and datasets to validate the superiority of the proposed method. In addition, we perform various ablation studies to demonstrate the generalization capability of the proposed method.
Figure 1: Visualization of the feature similarity. We show a feature similarity map using the features of the bicycle wheels area (red box) as the reference feature. In conventional methods, high feature similarity is observed with other bicycle wheels (yellow and blue boxes), whereas in the proposed method, feature similarity is notably reduced. (a) Clean image, (b) PGD [36] (c) SegPGD [17], (d) CosPGD [1], and (e) FSPGD (Ours).
2 Preliminaries
Given a source model F with parameters $\theta$ and a clean image x with ground-truth image y, the goal of attacker is to generate an adversarial example $\textrm{x}^{adv}$ that is indistinguishable from clean image x (i.e. $||\textrm{x}^{adv}-\textrm{x}||_{p}\leq\epsilon$) but can fool the source model $F(\textrm{x}^{adv};\theta)\neq F(\textrm{x};\theta)=\textrm{y}$. Here, $\epsilon$ indicates the perturbation budget, and |||||$\cdot$||| | ⋅ | | means the $l_{p}$ norm distance. In this paper, we set p as $\infty$ following conventional methods [17, 26, 53, 1, 5, 4]. To generate an adversarial example, the attacker typically maximizes the objective function which is defined as follows:
$\textrm{x}^{adv}=\underset{||\textrm{x}^{adv}-\textrm{x}||_{p}\leq\epsilon}{\mathrm{argmax}}L(\textrm{x}^{adv},\textrm{y};\theta),$ (1)
where L is the objective function defined by the user. For instance, in[15], $\textrm{x}^{adv}$ is generated in an intuitive manner as follows:
$\textrm{x}^{adv}=\textrm{x}+\epsilon\cdot\textrm{sign}(\nabla_{\textrm{x}}L(\textrm{x},\textrm{y};\theta)).$ (2)
This approach could efficiently produce adversarial examples but show poor attack performance. In [36], they introduce an iterative attack method, called projected gradient descent (PGD), which updates the adversarial example incrementally by adding small perturbations with a step size $\alpha$, which is expressed as
$\textrm{x}^{adv}_{t}=\textrm{x}^{adv}_{t-1}+\alpha\cdot\textrm{sign}(\nabla_{\textrm{x}^{adv}_{t}}L(\textrm{x}^{adv}_{t},\textrm{y};\theta)).$ (3)
Since PGD method shows better performance than single-step method defined in Eq. 2, following the previous papers [17, 1, 53, 3, 36, 38, 43, 47, 24], we employ the PGD as the baseline of the proposed method.
Recently, various adversarial attack methods [17, 26, 53, 1, 5, 4, 22] specialized for semantic segmentation have been introduced. For instance, Guo et al. [17] enhanced the existing projected gradient descent (PGD) method [36], originally developed for image classification, and demonstrated the effectiveness of the iterative attack strategy in semantic segmentation. Jia et al. [26] tried to further improve the transferability of the method introduced in [17] by designing a novel two-stage attack process. In [5], they proposed a new attack method by theoretically analyzing the limitation of the existing attack process, while Chen et al. [4] introduced a method to enhance attack transferability using an ensemble model. These methods show strong performance in the source model, but they have not yet fully overcome the challenges of transferability. More detailed explanations of related works are provided in the supplementary material.
3 Proposed Method
3.1 Motivation
We investigate the causes of weak transferability in conventional methods and identify the following issues. Conventional segmentation attacks [17, 26, 53, 1, 5, 4] typically aim to disrupt output predictions, similar to image classification attacks  [37, 10, 2, 16]. However, segmentation attacks differ fundamentally from image classification attacks. In image classification, an input image usually contains a single object representing one class. In semantic segmentation, however, the input image can contain multiple objects from different classes or multiple instances of the same class (e.g., multiple people). Traditional classification attack methods, developed under the assumption of a single object class, do not need to consider spatial relationships or contextual information. In contrast, segmentation attack methods must account for spatial relationships among objects within the input image. The most intuitive approach to disrupting spatial relationships is to generate an adversarial image where objects of the same class display dissimilar features, making correct predictions challenging.
Figure 2: Overall framework of FSPGD. FSPGD employs a loss function with two components: external and internal feature similarity loss. The external feature similarity loss measures similarity between intermediate-level features of the clean image and adversarial example, whereas the internal feature similarity loss compares intermediate-level feature similarity among similar objects within adversarial example.
To validate our hypothesis, we conducted experiments to visualize feature similarity in the intermediate layer, as depicted in Fig. 1. Using the feature vector of the bicycle wheel region (red box) as a reference feature, we generated a map comparing feature similarity with other areas, using DeepLabV3-ResNet50 as the source model and DeepLabV3-ResNet101 as the target model. As shown in Fig. 1(a), the clean image reveals that the reference feature is similar to those of other bicycle wheel regions (yellow and blue boxes), indicating that the network generates similar features for objects with the same class, even when they are spatially separated. Despite the attack, as shown in Figs. 1(b), (c), and (d), conventional methods still produce similar features in the target model. In other words, objects with the same class continue to exhibit similar features, leading to weak attack performance (producing predictions nearly identical to those for the clean image); these results show the low transferability in conventional methods. In contrast, the proposed method performs the attack by accounting for spatial relationships, resulting in feature dissimilarity between wheel regions (red, yellow, and blue boxes). Consequently, the proposed method achieves better attack performance and demonstrates superior transferability compared to conventional methods.
3.2 Methodology
Algorithm 1 Algorithm of FSPGD
Input: Clean image $\mathrm{x}$; clean image feature map $\textit{f}_{x}(\cdot)$; adversarial example feature map $\textit{f}_{a}(\cdot)$; attack iterations $T$; the maximum magnitude of adversarial perturbation $\epsilon$; step size $\alpha$; $\phi^{\epsilon}(\cdot)$ is a function that clips output into the range $[\mathrm{x}-\epsilon,\mathrm{x}+\epsilon]$; $\mathcal{U}(-\epsilon,\epsilon)$ is a function that initializes random noise into the range $[-\epsilon,\epsilon]$.
Output: The adversarial example $\mathrm{x}_{T}^{adv}$
1:Initialize $\mathrm{x}^{adv}_{0}=\mathrm{x}+\mathcal{U}(-\epsilon,\epsilon)$
2:for $t\leftarrow 0\ to\ T-1$ do
3:     $\lambda_{t}\leftarrow t/T$
4:     Calculate $L_{ex}$ by using Eq.(4)
5:     Calculate $L_{in}$ by using Eq.(8)
6:     $L=\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$
7:     Calculate the gradient of $L$ with respect to $\textrm{x}^{adv}_{t}$
8:     Update $\textrm{x}^{adv}_{t+1}$
$\textrm{x}^{adv}_{t+1}\leftarrow\textrm{x}^{adv}_{t}+\alpha\cdot sign(\triangledown_{\textrm{x}^{adv}_{t}}L)$
9:     Clamp on $\epsilon$-ball of clean image
$\textrm{x}^{adv}_{t+1}\leftarrow\phi^{\epsilon}(\textrm{x}^{adv}_{t+1})$
10:end for
In the proposed method, we build L function using the intermediate layer features $f\in\mathbb{R}^{c\times N}$, where c and N represent the number of channels and pixels of the feature map, respectively. In the remainder of this paper, we denote by $f_{x}\in\mathbb{R}^{c\times N}$ and $f_{a}\in\mathbb{R}^{c\times N}$, where the intermediate feature maps extracted from x and $\textrm{x}^{adv}_{t}$, respectively. Here, to successfully perform an attack on the source model, $f_{x}$ and $f_{a}$ should be as dissimilar as possible. Additionally, to ensure that similar objects exhibit different features in the intermediate layer of target models, similar objects within $f_{a}$ should have dissimilar vectors.
Based on this hypothesis, we design our framework as illustrated in Fig. 2. The proposed method consists of two different loss functions, i.e. $L_{ex}$ and $L_{in}$, which represent the external-feature similarity loss and internal-feature similarity loss, respectively. Specifically, $L_{ex}$ is a loss function designed to minimize the similarity between $f_{x}$ and $f_{a}$, aiming to successfully perform an attack on the source model. To achieve this, we design the loss function to reduce cosine similarity between feature vectors of each pixel in $f_{x}$ and $f_{a}$, which is formulated as follows:
$L_{ex}=\frac{1}{N}\sum\limits_{i=1}^{N}\Big{(}\frac{f_{x}(i)}{|f_{x}(i)|}\Big{)}^{\textrm{T}}\frac{f_{a}(i)}{|f_{a}(i)|},$ (4)
where i indicates the pixel location. This loss function is intuitive and simple, yet exhibits outstanding performance in semantic segmentation attacks.
Table 1: Attack performance comparison on Pascal VOC 2012 in terms of mIoU. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup
[…]
Source Models Attack Method Source Model PSPRes101 DV3Res101 FCNVGG16
Clean Images 80.22/80.18 78.39 82.88 59.80
PSPRes50 PGD [36] 7.72 54.73 59.41 45.70
SegPGD [17] 5.41 54.10 58.95 45.43
CosPGD [1] 1.84 56.63 64.37 45.99
DAG [53] 65.82 62.67 66.22 38.91
NI [32] 7.71 33.49 38.52 32.94
DI [54] 6.41 32.00 35.25 37.34
TI [11] 18.28 64.50 69.60 36.80
FSPGD (Ours) 3.39 22.24 16.84 19.75
DV3Res50 PGD [36] 9.74 52.96 56.35 46.39
SegPGD [17] 7.26 52.05 56.50 46.23
CosPGD [1] 1.67 56.82 61.36 45.94
DAG [53] 66.78 62.12 66.84 38.77
NI [32] 9.89 33.86 36.85 34.92
DI [54] 7.35 31.93 32.93 38.30
TI [11] 19.34 64.99 69.80 37.65
FSPGD(Ours) 3.44 21.89 16.57 19.36
On the other hand, $L_{in}$ is designed to generate dissimilar features for similar objects within the image, addressing the issues discussed in Sec. 3.1. We first measure the similarity of $f_{a}$ between each pixel and all other pixels by constructing the Gram matrix $\textrm{S}\in\mathbb{R}^{N\times N}$ as follows:
$\textrm{S}(p,q)=\Big{(}\frac{f_{a}(p)}{|f_{a}(p)|}\Big{)}^{\textrm{T}}\frac{f_{a}(q)}{|f_{a}(q)|},$ (5)
where $p=1,2,...,N$ and $q=1,2,...,N$. Note that our goal is to perform the attack only on pixels corresponding to regions with similar objects, rather than on all pixels. That means, we have to identify the locations of similar objects within the clean image based on the observation that similar objects have similar features. To this end, we design a mask matrix $\textrm{M}\in\mathbb{R}^{N\times N}$ for selecting pixels containing similar objects, where M is defined as
$\textrm{M}(p,q)=\Big{(}\frac{f_{x}(p)}{|f_{x}(p)|}\Big{)}^{\textrm{T}}\frac{f_{x}(q)}{|f_{x}(q)|}.$ (6)
Here, we build M using the $f_{x}$ instead of $f_{a}$ since $f_{x}$ always retains the same features, regardless of the progression of the attack. Note that when $f_{x}(p)$ and $f_{x}(q)$ have similar features due to similar objects, $\textrm{M}(p,q)$ would have a high value; that means p-th and q-th pixels have strong spatial relationships. Indeed, since M contains numerous components (e.g. when N is 1,024, i.e. $32\times 32$ resolution, M has approximately 1 million components), it is challenging to cover all pixels correlations. Thus, we simplify M and select specific pixels by performing binarization as follows:
$\textrm{M}_{B}(p,q)=\begin{cases}1,&\text{if $\textrm{M}(p,q)>\tau$}\\ 0,&\text{otherwise}\end{cases}$ (7)
Table 2: Attack performance comparison on Cityscapes in terms of mIoU. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup
[…]
Source Models Attack Source PSP DV3 PSP DV3 Mask2Former
Method Model Res50 Res50 Res101 Res101 Swin-S
Clean Images 60.58 64.62 65.65 65.90 67.16 68.24
SegFormer PGD [36] 1.06 29.94 36.07 31.99 38.25 48.43
MiT-B0 SegPGD [17] 0.38 28.45 34.56 29.28 36.38 49.54
CosPGD [1] 0.00 29.98 35.92 32.19 37.72 51.51
DAG [53] 50.92 20.84 33.73 32.71 28.77 55.21
NI [32] 2.06 30.27 37.63 30.95 38.24 43.75
DI [54] 9.13 41.92 45.85 43.10 48.06 46.78
TI [11] 7.66 50.60 52.77 52.25 55.88 55.59
FSPGD (Ours) 1.33 10.09 14.57 21.16 22.06 39.92
Source Models Attack Source PSP DV3 PSP DV3 SegFormer
Method Model Res50 Res50 Res101 Res101 MiT-B0
Clean Images 68.24 64.62 65.65 65.90 67.16 60.58
Mask2Former PGD [36] 0.45 39.41 45.25 42.15 48.35 49.30
Swin-S SegPGD [17] 0.30 39.97 45.07 42.29 48.96 49.40
CosPGD [1] 0.17 39.56 45.23 42.36 47.43 49.37
DAG [53] 65.59 30.69 42.06 32.76 39.42 54.23
NI [32] 0.17 42.76 49.41 45.06 50.00 45.87
DI [54] 3.53 50.34 53.67 53.16 56.59 50.85
TI [11] 0.85 56.81 59.74 59.95 62.69 59.74
FSPGD (Ours) 2.20 15.57 18.00 24.29 25.96 36.87
where $\tau$ is an user-defined threshold value. By using Eqs. 5 and 7, we define $L_{in}$ as follows:
$L_{in}=\frac{1}{2}\frac{1}{{K}}\sum\limits_{p=1}^{N}\sum\limits_{q=1}^{N}\textrm{M}_{B}(p,q)\otimes\textrm{S}(p,q),$ (8)
where $\otimes$ indicates element-wise multiplication operation and K is the number of elements with a value of 1 in the $\textrm{M}_{B}$ matrix (i.e. $K=\sum_{p}\sum_{q}\textrm{M}_{B}(p,q)$). Since both $\textrm{M}_{B}$ and S are symmetric Gram matrices, we divided by two to avoid double-counting values (i.e. 1/2 in Eq. 8).
By combining Eqs. 4 and 8, we make our objective function L as follows:
$L=\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in},$ (9)
where $\lambda_{t}$ is a value that controls the balance between $L_{ex}$ and $L_{in}$. Through extensive experiments, we found that it is beneficial to use $L_{in}$ in the early stages of attack iterations to reduce feature similarity between objects of the same class, and to apply $L_{ex}$ in the later stages to reduce the similarity between $f_{x}$ and $f_{a}$. Based on these observations, we define $\lambda_{t}=t/T$. Extensive experiments on the value of $\lambda_{t}$ are provided in the ablation study and supplementary material. We summarize the algorithm of the proposed method in Algorithm 1.
4 Experiment
Figure 3: Visualization of experimental results. DV3Res50 is used as the source model and images of first column are clean images and adversarial examples generated by PGD [36], SegPGD [17], CosPGD [1], and FSPGD (Ours). second column is ground truth of input images. And other columns are predictions of target models.
4.1 Experimental Setup
Datasets. We use two popular semantic segmentation datasets in our experiments: PASCAL VOC 2012 [12], Cityscapes [9]. The VOC dataset includes 20 object classes and one background class, containing 1,464 images for training and 1,499 for validation. Following the standard protocol [20], the training set is expanded to 10,582 images. The Cityscapes dataset, focused on urban scene understanding, comprises 19 categories with high-quality pixel-level annotations, including 2,975 images for training and 500 for validation. In our experiments, attack performance is evaluated using the validation set of each dataset.
Models. In this paper, we employ popular semantic segmentation models, i.e. PSPNet-ResNet50, DeepLabv3-ResNet50 [6], SegFormer-MiT B0 [55], and Mask2Former-Swin S [8] as our source and target models, with FCN-VGG16 [34] additionally used as target model. We conduct cross-validation by alternating source and target models to demonstrate the transferability of the proposed method. For instance, when PSPNet-ResNet50 is used as the source model, we measure attack performance on DeepLabv3-ResNet101, PSPNet-ResNet101, FCN-VGG16.
Parameters. Each comparison experiment follows the $l_{\infty}$-norm, setting the maximum perturbation value $\epsilon$ to 8/255. The step size $\alpha$ is set to 2/255 and the total iteration $T$ is set to 20. The proposed method has a user parameter $\tau$ which acts the threshold value in Eq. 7. In our experiments, we set $\tau$ value as cos($\pi/3$). The reason we set the threshold value as a cosine value is as follows: since $\textrm{M}(p,q)$ is calculated through the inner product of two vectors with a magnitude of 1, its value represents the cosine of the angle $\theta$ between two vectors. Therefore, we choose the threshold value based on the cosine value.
Metrics. To assess the adversarial robustness of segmentation models, we use the standard metric, mean Intersection over Union (mIoU). Lower mIoU indicate greater attack performance. We report mIoU (%) scores for both clean images and adversarial examples.
4.2 Experimental Results
We first compare the attack performance on conventional methods [1, 17, 36, 53, 32, 54, 11] with the proposed method. The experimental results are summarized in Tables 1 and  2. PSPResX and DV3ResX indicate the PSPNet [58] and DeepLabV3 [6] with ResNet50 [21] (or ResNet101 [21]) encoder, respectively. In the Cityscapes, SegFormer [55] with a MiT-B0 [55] encoder and Mask2Former [8] with a Swin-S [33] encoder are used as transformer-based source models. The proposed method shows high attack performance on the source model compared to conventional methods, excluding CosPGD [1] which is designed for white-box attack. To evaluate transferability, we measure mIoU on various target models. In this study, we select target models such that the encoders (e.g. ResNet50 and ResNet101) do not overlap between source and target models. As shown in Tables 1 and 2, the proposed method exhibits significantly superior transferability compared to conventional methods. In particular, it shows strong attack performance not only on target models using ResNet-based encoders but also on substantially different models based on transformer in Table 2. These results indicate that the proposed method is better suited for real-world scenarios compared to traditional methods. Due to page limitations, we compare the performance of only a few source models in Tables 1 and  2. Additional experimental results comparing a wider range of conventional methods are described in the supplementary material.
For qualitative evaluation, we visualize adversarial examples along with their corresponding prediction results. In our experiments, we set DeepLabV3-ResNet50 as the source model. PSPNet with Resnet50 (and Resnet101)and DeepLabV3 with Resnet50 (and Resent101) set as the target model. As shown in Fig. 3, prediction results of conventional methods are similar to the results on clean images, indicating weak transferability. In contrast, the proposed method successfully attacks target models, demonstrating strong transferability. Based on these results, we conclude that the proposed method achieves the state-of-the-art transferability performance. Additional images of attack results are provided in the supplementary material.
4.3 Ablation Studies
Figure 4: mIoU performance across different loss terms. (S) and (T) indicate the source and target models, respectively.
The proposed method incorporates a user-defined variable $\tau$ for binarizing M. To determine the optimal $\tau$ value, we conduct ablation studies on Pascal VOC 2012 dataset. To select the $\tau$ value that maximizes transferability, we conduct experiments with all other variables fixed by setting $\tau$ to $cos(\pi/3)$, $cos(\pi/4)$, and $cos(\pi/6)$ and Fig. 4 presents the results. Since the average mIoU value was the lowest when $\tau$ was set to $cos(\pi/3)$, we selected $cos(\pi/3)$ in our study. As shown in Table 1 and Fig. 4, the proposed method outperforms existing methods, regardless of the $\tau$ value. Therefore, we believe that, despite having a user-defined parameter, the proposed method offers the advantage of superior performance compared to existing methods. The proposed loss function consists of two components: $L_{ex}$ and $L_{in}$. To evaluate the effect of each loss term, we conduct an ablation study with five different configurations. First, we consider using only the external loss term $L_{ex}$, and second, using only the internal loss term $L_{in}$. Third, we explore a fixed weight combination of the two losses, where $L_{ex}$ + $\lambda_{t}L_{in}$ is used with a constant $\lambda$ (e.g. 0.1, 0.5, 1.0). Finally, we employ an adaptive weighting strategy in which both losses are dynamically adjusted using $\lambda$ at each iteration, following the formulation $\lambda_{t}L_{ex}$ + $(\lambda_{t}-1)L_{in}$. As shown in Fig. 5, we observe that the experiment using the dynamic lambda strategy achieved better attack performance compared to other loss combinations. Hence, we employ the dynamic lambda strategy.
Furthermore, we conduct ablation studies to determine the most effective feature extraction layers in various source models. Trough conducting extensive experiments, we determine the optimal layers are Layer 2 of Conv3_x in ResNet-50, Layer 1 of Transformer block 1 in MiT-B0, and Layer 1 of Stage 2 in Swin-S. Detailed results are provided in the Supplemental Material.
5 Limitations
The proposed method demonstrates superior transferability compared to existing methods. However, a drawback of the proposed method is the presence of the user-defined parameter $\tau$ and loss balance parameter $\lambda_{t}$. While the ablation study illustrates performance variations according to different $\tau$ values, there would be better $\tau$ which leads higher attack performance. Additionally, we observe that attack performance varies depending on how the two loss terms, $L_{ex}$ and $L_{in}$, are adjusted through the $\lambda_{t}$ value. Ideally, the $\tau$ and $\lambda_{t}$ values should be determined automatically by taking into account the characteristics of the input image, the source model, and feature distributions. In future research, we plan to investigate techniques for automatically selecting optimal $\tau$ and $\lambda_{t}$ values.
Figure 5: mIoU performance across different loss terms. (S) and (T) indicate the source and target models, respectively.
6 Conclusion
In this paper, we identify key limitations in existing segmentation attack methods and conduct an in-depth analysis of the underlying causes. Based on these observations, we develop and introduce a novel segmentation attack method, called Feature Similarity Projected Gradient Descent (FSPGD), specifically designed to enhance both attack performance and transferability. The proposed FSPGD method demonstrates notable improvements over conventional methods, not only in terms of attack efficacy but also in transferability across different model architectures. Future work will aim to further optimize the parameter settings of FSPGD to enhance its robustness and adaptability across various model configurations. Additionally, we plan to explore a more automated approach for parameter optimization, which would allow the method to achieve optimal results efficiently across a diverse set of models, thus broadening its applicability in real-world scenarios.
Acknowledgement. This research was supported by the MSIT (Ministry of Science and ICT), Korea under the ITRC(Information Technology Research Center) support program (IITP-2024-RS-2023-00258971) supervised by the IITP (Institute for Information & Communications Technology Planning & Evaluation).
[…]
Supplementary Material
A Visualization of Feature Similarity
To further validate the motivation described in Sec. 3.1 , we performed visualizations on a broader variety of images. Figs. 1 and  2 present experimental results on the Pascal VOC 2012 dataset, while Figs. 3 and Figs. 4 show results on the Cityscapes dataset. As seen in the figures, conventional methods maintain the similarity of features within the same class even after performing an attack, leading to poor attack performance on new target models. In contrast, the proposed method reduces feature similarity and exhibits superior attack performance compared to conventional methods.
Figure 1: Visualization of the feature similarity on Pascal VOC 2012 dataset. Red boxes indicate the reference features, while yellow and blue boxes represent regions belonging to the same class as the red boxes. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 is used as target model. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 2: Visualization of the feature similarity on Pascal VOC 2012 dataset. Red boxes indicate the reference features, while yellow and blue boxes represent regions belonging to the same class as the red boxes. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 is used as target model. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 3: Visualization of the feature similarity on Cityscapes dataset. Red boxes indicate the reference features, while yellow and blue boxes represent regions belonging to the same class as the red boxes. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 is used as target model. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 4: Visualization of the feature similarity on Cityscapes dataset. Red boxes indicate the reference features, while yellow and blue boxes represent regions belonging to the same class as the red boxes. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 is used as target model. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
B Extended Experimental Results
To further prove the superiority of the proposed method, we conducted comparative experiments with various conventional methods [36, 17, 1, 53, 32, 54, 11]. To evaluate the transferability of the attack methods, we designed the experiments with non-overlapping encoders for the source model and target model. As shown in Table 1 and  2, the proposed method achieves the best performance among black-box attack methods on the source model (CosPGD [1] is a white-box attack method) and shows superior attack performance on target models compared to existing methods.
Table 1: Attack performance comparison on Pascal VOC 2012 in terms of mIoU. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup
[…]
Source Models Attack Method Source Model PSPRes101 DV3Res101 FCNVGG16
Clean Images 80.22/80.18 78.39 82.88 59.80
PSPRes50 PGD [36] 7.72 54.73 59.41 45.70
SegPGD [17] 5.41 54.10 58.95 45.43
CosPGD [1] 1.84 56.63 64.37 45.99
DAG [53] 65.82 62.67 66.22 38.91
NI [32] 7.71 33.49 38.52 32.94
DI [54] 6.41 32.00 35.25 37.34
TI [11] 18.28 64.50 69.60 36.80
FSPGD (Ours) 3.39 22.24 16.84 19.75
DV3Res50 PGD [36] 9.74 52.96 56.35 46.39
SegPGD [17] 7.26 52.05 56.50 46.23
CosPGD [1] 1.67 56.82 61.36 45.94
DAG [53] 66.78 62.12 66.84 38.77
NI [32] 9.89 33.86 36.85 34.92
DI [54] 7.35 31.93 32.93 38.30
TI [11] 19.34 64.99 69.80 37.65
FSPGD(Ours) 3.44 21.89 16.57 19.36
Source Models Attack Method Source Model PSPRes50 DV3Res50 FCNVGG16
Clean Images 78.39/82.88 80.22 80.18 59.80
PSPRes101 PGD [36] 10.13 55.39 55.39 47.25
SegPGD [17] 7.31 53.56 54.03 46.26
CosPGD [1] 2.87 57.74 58.50 47.05
DAG [53] 63.36 66.28 66.06 39.10
NI [32] 10.22 33.50 34.12 34.41
DI [54] 7.21 29.00 30.58 39.24
TI [11] 22.23 64.64 64.95 37.29
FSPGD(Ours) 2.99 12.48 13.54 21.30
DV3Res101 PGD [36] 9.75 59.36 55.54 47.48
SegPGD [17] 7.18 54.47 53.96 46.53
CosPGD [1] 2.73 58.83 58.54 47.25
DAG [53] 67.55 67.09 67.58 39.48
NI [32] 9.49 36.41 34.75 35.62
DI [54] 7.64 34.87 34.11 40.99
TI [11] 27.16 65.79 65.13 37.98
FSPGD(Ours) 3.28 11.42 13.45 21.49
Table 2: Attack performance comparison on Cityscapes in terms of mIoU. Lower mIoU means better performance and bold numbers denote the best mIoU values for each experimental setup
[…]
Source Models Attack Method Source PSP DV3 Segformer Maskformer
Model Res101 Res101 MiT-B0 Swin-S
Clean Images 64.62 / 65.90 65.65 67.16 60.58 68.24
PSP PGD [36] 1.83 18.80 19.35 48.92 59.66
Res50 SegPGD [17] 1.38 18.26 19.34 49.83 60.41
CosPGD [1] 0.07 24.90 26.65 50.31 60.53
DAG [53] 23.52 36.76 33.47 50.24 60.64
NI [32] 1.62 15.07 17.07 44.43 50.09
DI [54] 1.92 17.60 21.57 52.12 54.81
TI [11] 1.64 28.39 34.07 51.91 58.70
FSPGD (Ours) 0.93 5.12 3.29 41.30 47.30
DV3 PGD [36] 2.00 22.19 22.06 50.28 60.64
Res50 SegPGD [17] 0.96 22.20 22.51 50.59 60.24
CosPGD [1] 0.01 25.43 27.22 50.48 59.86
DAG [53] 36.54 39.01 35.98 51.59 60.19
NI [32] 1.55 16.65 18.26 45.76 49.89
DI [54] 2.32 19.87 23.61 52.63 55.32
TI [11] 1.48 31.93 35.45 52.77 59.56
FSPGD (Ours) 1.27 6.09 3.78 40.74 47.30
Source Models Attack Method Source PSP DV3 Segformer Maskformer
Model Res50 Res50 MiT-B0 Swin-S
Clean Images 65.65 / 67.16 64.62 65.90 60.58 68.24
PSP PGD [36] 1.80 9.71 12.80 48.83 59.57
Res101 SegPGD [17] 0.90 10.64 12.85 60.41 59.48
CosPGD [1] 0.02 14.02 16.41 50.75 61.01
DAG [53] 35.74 23.56 33.92 51.65 61.42
NI [32] 1.65 8.35 10.01 42.51 46.90
DI [54] 2.18 16.94 19.39 50.57 53.31
TI [11] 1.73 25.15 29.84 50.95 57.30
FSPGD (Ours) 2.29 5.96 7.42 36.63 36.91
DV3 PGD [36] 1.74 15.20 16.54 49.92 60.50
Res101 SegPGD [17] 0.63 17.29 18.04 50.18 60.11
CosPGD [1] 0.01 18.83 19.60 50.44 59.91
DAG [53] 36.68 26.70 36.68 52.25 61.29
NI [32] 1.94 14.15 14.91 44.14 48.74
DI [54] 3.99 22.41 23.87 50.91 53.96
TI [11] 2.63 29.58 32.17 51.90 57.19
FSPGD (Ours) 2.03 2.48 3.25 39.82 47.04
C Additional Examples for Qualitative Evaluation
Figure 5: Visualization of clean image, attacked images, and output predictions on Pascal VOC 2012. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row), and PSPNet-Res101 (third row) are used as target models. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 6: Visualization of clean image, attacked images, and output predictions on Pascal VOC 2012. Deeplabv3-Res50 is used as the source model and Deeplabv3-Res101 (second row), and PSPNet-Res101 (third row) are used as target models. are used as target models. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 7: Visualization of clean image, attacked images, and output predictions on Cityscapes. Deeplabv3-Res101 is used as the source model and Deeplabv3-Res50 (second row), Segformer-MiT B0 (third row), and Mask2former-SwinS (fourth row) are used as target models. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
Figure 8: Visualization of clean image, attacked images, and output predictions on Cityscapes. Deeplabv3-Res101 is used as the source model and Deeplabv3-Res50 (second row), Segformer-MiT B0 (third row), and Mask2former-SwinS (fourth row) are used as target models. (a) Clean image, (b) PGD [36], (c) SegPGD [17], (d) CosPGD [1], (e) FSPGD (Ours).
D Detailed Experimental Results for Ablation Studies
This section presents the quantified experimental results in ablation studies discussed in Sec. 4.3 and provides a more detailed explanation of these results. Additionally, it elaborates on ablation study findings that were not included in the main text due to space constraints.
D.1 Performance comparison based on $\tau$ value
The proposed method includes a user-defined parameter, $\tau$, which is used to build the mask $\textrm{M}_{B}$. Since the $\tau$ value affects the attack performance, we conducted extensive experiments to compare the results. As shown in Table 3 and Table 4, the attack performance varies slightly depending on the $\tau$ value. Notably, although performance fluctuates with different $\tau$ values, it consistently outperforms conventional techniques shown in Table 1 and 2 in main paper. We calculated the average performance for each $\tau$ value and selected cos($\pi/3$) as the optimal value, as it achieved the highest average performance.
Table 3: Attack performance comparison in Pascal VOC 2012 dataset across different $\tau$ values. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
  Source Models $\tau$ Target Models
Source Model PSPRes101 DVRes101 FCNVGG16
PSPRes50 $\pi/6$ 3.37 28.49 22.05 21.39
$\pi/4$ 3.40 24.53 17.92 20.44
$\pi/3$ 3.39 22.24 16.84 19.75
DV3Res50 $\pi/6$ 3.46 27.52 22.01 20.98
$\pi/4$ 3.45 23.85 17.77 20.16
$\pi/3$ 3.44 21.89 16.57 19.36
Source Models $\tau$ Target Models
Source Model PSPRes50 DVRes50 FCNVGG16
PSPRes101 $\pi/6$ 3.01 20.89 20.30 24.10
$\pi/4$ 3.04 11.77 12.30 21.55
$\pi/3$ 2.99 12.48 13.54 21.30
DV3Res101 $\pi/6$ 3.29 21.43 20.69 24.52
$\pi/4$ 3.25 11.37 11.95 21.57
$\pi/3$ 3.28 11.42 13.45 21.49
Table 4: Attack performance comparison in Cityscapes dataset across different $\tau$ values. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
[…]
Source Models $\tau$ Source PSP DV3 PSP DV3 Mask2former
Model Res50 Res50 Res101 Res101 Swin-S
Segformer $\pi/6$ 1.70 9.98 16.38 25.98 25.08 43.74
MiT-B0 $\pi/4$ 1.36 11.19 16.59 24.76 24.81 42.97
$\pi/3$ 1.33 10.09 14.57 21.16 22.06 39.92
Source Models $\tau$ Source PSP DV3 PSP DV3 Segformer
Model Res50 Res50 Res101 Res101 MiT-B0
Mask2former $\pi/6$ 4.15 16.83 20.47 26.13 28.58 38.84
Swin-S $\pi/4$ 2.94 16.62 19.61 24.64 26.78 37.80
$\pi/3$ 2.20 15.57 18.00 24.29 25.96 36.87
D.2 Performance comparison based on $\lambda$ value
The proposed loss function consists of two loss terms, $L_{ex}$ and $L_{in}$. Here, we provide a detailed numerical explanation of the experimental results, along with additional results for loss term combinations. Table 3 summarizes the experimental results on the Pascal VOC 2012 and Cityscapes dataset. As shown in Table 5 and 6, the performance of the proposed method varies depending on how the two loss terms are combined. As discussed in the main text, simply adding the two loss terms can result in a compromise, leading to lower performance compared to using $L_{ex}$ alone. To investigate this performance degradation, we conducted experiments with different ratios, such as $L_{ex}+0.5L_{in}$ and $L_{ex}+0.1L_{in}$. The results, as summarized in the Table 5 and 6, show that performance varies depending on the source model; for instance, when PSPNet-Res50 is the source model, performance was lower compared to using $L_{ex}$ alone, but when DeepLabv3-Res50 was used, performance improved. To address this issue of performance variation across source models, we proposed a dynamic $\lambda_{t}$ that adjusts with t, and this method demonstrated the best performance overall.
Table 5: Attack performance comparison across different loss combinations. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
  Target Models
Source Models $\lambda$ Source Model PSPRes101 DVRes101 FCNVGG16
PSPRes50 $L_{ex}$ 3.37 24.81 18.13 20.01
$L_{in}$ 4.06 60.96 61.92 37.09
$L_{ex}+L_{in}$ 3.40 25.78 19.11 20.98
$L_{ex}+0.5L_{in}$ 3.41 25.07 18.51 20.44
$L_{ex}+0.1L_{in}$ 3.37 24.93 18.16 20.13
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 3.39 22.24 16.84 19.75
DV3Res50 $L_{ex}$ 3.47 25.19 18.93 19.78
$L_{in}$ 4.01 58.42 58.87 36.90
$L_{ex}+L_{in}$ 3.45 24.69 19.35 20.54
$L_{ex}+0.5L_{in}$ 3.45 24.76 18.73 20.08
$L_{ex}+0.1L_{in}$ 3.45 24.86 18.64 19.75
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 3.44 21.89 16.57 19.36
Source Models $\lambda$ Source Model PSPRes50 DVRes50 FCNVGG16
PSPRes101 $L_{ex}$ 3.13 14.36 14.68 21.03
$L_{in}$ 4.75 45.67 48.64 37.59
$L_{ex}+L_{in}$ 3.04 12.97 14.57 21.41
$L_{ex}+0.5L_{in}$ 3.06 13.17 14.22 21.08
$L_{ex}+0.1L_{in}$ 3.12 13.83 14.48 21.02
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 2.99 12.48 13.54 21.30
DV3Res101 $L_{ex}$ 3.32 12.60 13.44 20.57
$L_{in}$ 17.54 65.77 66.57 39.99
$L_{ex}+L_{in}$ 3.71 32.42 30.21 23.37
$L_{ex}+0.5L_{in}$ 3.61 29.84 27.75 22.20
$L_{ex}+0.1L_{in}$ 3.57 27.86 25.17 21.12
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 3.28 11.42 13.45 21.49
Table 6: Attack performance comparison across different loss combinations. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
  Target Models
Source Models $\lambda$ Source PSP DV PSP DV Mask2former
Model Res50 Res50 Res101 Res101 Swin-S
Segformer $L_{ex}$ 60.20 55.94 60.08 60.20 64.11 64.62
MiT-B0 $L_{in}$ 21.73 30.56 28.59 36.75 37.91 47.93
$L_{ex}+L_{in}$ 56.43 52.85 56.40 57.59 60.04 63.53
$L_{ex}+0.5L_{in}$ 58.46 54.55 58.78 59.34 62.27 64.17
$L_{ex}+0.1L_{in}$ 59.92 55.61 59.80 60.47 63.78 64.49
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 1.33 10.09 14.57 21.16 22.06 39.92
Source Models $\lambda$ Source PSP DV PSP DV Segformer
Model Res50 Res50 Res101 Res101 MiT-B0
Maskformer $L_{ex}$ 67.05 58.61 61.84 61.77 64.21 58.58
Swin-S $L_{in}$ 24.05 44.91 44.76 44.77 50.27 51.84
$L_{ex}+L_{in}$ 64.58 55.94 59.34 59.00 62.06 57.42
$L_{ex}+0.5L_{in}$ 65.99 57.40 60.67 60.79 63.28 57.97
$L_{ex}+0.1L_{in}$ 67.43 58.38 61.59 61.87 64.38 58.52
$\lambda_{t}L_{ex}+(1-\lambda_{t})L_{in}$ 2.20 15.57 18.00 24.29 25.96 36.87
D.3 Performance comparison based on layer location
Unlike conventional methods, the proposed method performs attacks by leveraging intermediate-layer features, making it the first approach to introduce intermediate-layer attacks in the field of semantic segmentation. As such, unlike intermediate-layer attack methods in image classification, there is no prior research on which layer is optimal for attacks in semantic segmentation. To address this, we conducted extensive experiments by attacking various layers of the encoder and summarized the results. Tables 7,  8, and  9 present the intermediate-layer attack performance for ResNet50, ResNet101 encoders, and transformer encoders, respectively. Attacking the later layers of the encoder (i.e., layer 4_2) results in strong performance on the source model but poor performance on the target models. In contrast, attacking the middle layers demonstrates reasonable attack performance on the source model while also achieving high transferability. Therefore, as the proposed method aims to enhance transferability, we chose to attack the middle layers.
Table 7: Attack performance results on source models using the ResNet50 encoder across different attack layers, evaluated on the Pascal VOC 2012 dataset. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
  Source Models Layer name Target Models
Source Model PSPRes101 DVRes101 FCNVGG16
PSPRes50 2_1 11.42 50.11 50.89 22.95
2_2 5.75 44.51 43.05 23.90
2_3 5.87 45.12 41.42 25.22
3_1 3.45 26.42 20.49 19.34
3_2 3.39 22.24 16.84 19.75
3_3 3.37 22.39 16.84 21.36
3_4 3.28 24.35 18.74 22.30
3_5 3.24 26.04 19.83 24.03
4_1 2.82 52.72 51.89 34.52
4_2 1.92 69.88 72.03 42.22
DV3Res50 2_1 8.11 48.15 47.67 22.10
2_2 4.45 41.36 38.61 22.42
2_3 4.74 41.45 38.47 23.86
3_1 3.47 26.37 20.33 18.81
3_2 3.44 21.89 16.57 19.36
3_3 3.39 22.19 17.47 20.93
3_4 3.35 24.19 19.22 22.22
3_5 3.26 24.99 19.52 23.77
4_1 2.38 50.66 51.72 34.40
4_2 2.36 67.71 69.74 41.08
Table 8: Attack performance results on source models using the ResNet101 encoder across different attack layers, evaluated on the Pascal VOC 2012 dataset. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
  Source Models Layer name Target Models
Source Model PSPRes50 DVRes50 FCNVGG16
PSPRes101 2_1 17.14 50.60 44.57 24.51
2_2 5.84 37.81 33.03 22.56
2_3 5.41 38.71 33.99 23.97
3_1 3.11 31.59 29.55 22.23
3_2 3.46 34.19 30.77 23.29
3_5 3.44 17.41 17.12 19.48
3_10 2.99 12.48 13.54 21.30
3_15 3.05 18.20 17.82 24.12
3_20 3.05 36.45 35.41 31.44
3_22 2.93 40.76 41.55 34.30
4_1 3.11 56.98 55.85 37.44
4_2 2.78 65.26 64.50 41.44
DV3Res101 2_1 17.78 51.02 44.56 24.57
2_2 7.40 37.30 32.94 22.94
2_3 6.38 41.47 34.65 24.49
3_1 3.36 32.32 31.56 22.26
3_2 3.57 33.69 32.74 23.79
3_5 3.46 17.33 17.68 19.85
3_10 3.28 11.42 13.45 21.49
3_15 3.35 18.55 19.01 25.20
3_20 3.38 38.31 39.72 33.36
3_22 3.25 43.07 45.80 35.58
4_1 3.17 57.74 56.16 38.35
4_2 1.49 63.18 62.93 40.99
Table 9: Attack performance results on source models using the transformer encoder across different attack layers, evaluated on the Cityscapes dataset. We measured mIoU scores and bold numbers indicate the best performance for each experimental setup.
  Target Models
Source Layer name Source PSP DV3 PSP DV3 Mask2former
Models Model Res50 Res50 Res101 Res101 Swin-S
Segformer patch_embeddings-0 43.55 25.45 27.83 30.83 33.77 51.55
MiT-B0 patch_embeddings-1 2.54 10.88 17.94 23.39 25.40 43.25
patch_embeddings-2 1.53 13.56 17.04 27.39 25.47 42.05
patch_embeddings-3 0.67 13.97 15.49 25.53 22.75 44.06
block-0-0 22.11 16.02 23.01 26.31 28.99 51.51
block-0-1 15.28 12.53 18.58 24.94 28.53 46.98
block-1-0 1.33 10.09 14.57 21.16 22.06 39.92
block-1-1 1.16 9.02 13.11 24.03 21.58 43.41
block-2-0 1.01 12.72 13.80 22.98 20.18 40.12
block-2-1 0.86 11.85 13.94 23.82 21.94 41.06
block-3-0 0.33 17.79 18.56 27.58 25.99 45.51
block-3-1 0.16 24.99 25.54 33.59 34.81 50.41
Source Layer name Source PSP DV3 PSP DV3 Segformer
Models Model Res50 Res50 Res101 Res101 MiT-B0
Mask2former embeddings 57.96 27.09 40.40 34.54 44.27 53.06
Swin-S layers-0-blocks-0 48.77 23.98 38.13 29.91 42.71 50.91
layers-0-blocks-1 37.73 16.41 33.17 24.86 36.57 46.63
layers-1-blocks-0 33.11 19.11 25.37 23.72 29.57 44.34
layers-1-blocks-1 17.77 13.65 25.81 19.64 28.03 41.63
layers-2-blocks-0 2.20 15.57 24.29 18.00 25.96 36.87
layers-2-blocks-17 0.79 30.47 36.66 31.20 38.85 45.15
layers-3-blocks-0 1.11 31.94 38.83 34.20 41.06 45.86
layers-3-blocks-1 1.41 33.83 40.14 36.43 43.11 46.49